hello i cant figure out this question all it saids is "graph" this is the eqation
x2+y2+4x-8y=16
thanks i would really appreciate if i can get help wit this

Found 2 solutions by mgmoeab, smartguy101:
Answer by mgmoeab(37)

** put all like- terms together.
x2+y2+4x-8y=16
x^2 + 4x + y^2 - 8y = 16
**complete the squares
(x^2 + 4x + 4) + (y^2 -8y + 16) = 16 + 4 +16
(x + 2)^ 2 + (y - 4)^2 = 4
** as you can see thats a circle of radius 2 and center ( -2, 4).
Hope that helps.

Answer by smartguy101(5)

put all like terms together.
x2+y2+4x-8y=16
x^2 + 4x + y^2 - 8y = 16
complete the squares
(x^2 + 4x + 4) + (y^2 -8y + 16) = 16 + 4 +16
(x + 2)^ 2 + (y - 4)^2 = 4 circle of radius 2 and center ( -2, 4)